OpenStudy (anonymous):

What is the missing statement in the proof? Scroll down to see the entire proof. ∠BDC≅∠ADB ∠BCA≅∠DCB ∠BAC≅∠BAD ∠DBC≅∠BAC

OpenStudy (anonymous):

? What is the missing statement in the proof? Scroll down to see the entire proof. ∠BDC≅∠ADB ∠BCA≅∠DCB ∠BAC≅∠BAD ∠DBC≅∠BAC Done Given: ΔABC with m∠ABC = 90° (view diagram) Prove: AB2 + BC2 = AC2 Statement Reason 1. Draw BD¯¯¯¯¯⊥AC¯¯¯¯¯ (view diagram). construction 2. ∠ABC≅∠BDC Angles with the same measure are congruent. 3. Reflexive Property of Congruence 4. ΔABC~ΔBDC AA criterion for similarity 5. BCDC=ACBC Corresponding sides of similar triangles are proportional. 6. BC2 = AC × DC cross multiplication 7. ∠ABC≅∠ADB Angles with the same measure are congruent. 8. ∠BAC≅∠DAB Reflexive Property of Congruence 9. ΔABC~ΔADB AA criterion for similarity 10.ABAD=ACAB Corresponding sides of similar triangles are proportional. 11. AB2 = AC × AD cross multiplication 12. AB2 + BC2 = AC × AD + AC × DC addition 13. AB2 + BC2 = AC(AD + DC) Distributive Property 14. AB2 + BC2 = AC × AC segment addition 15. AB2 + BC2 = AC2 multiplication

OpenStudy (anonymous):

@mr.chrzanowski

OpenStudy (anonymous):

yep he right